# Exploring Pintopia: Reflection Branes, Bordisms, and U-Dualities

### v1

@@ 1 Introduction
Similar considerations hold for the dynamics of multiple reflection branes. We find that for reflections associated with an even number of distinct internal directions, the resulting branes form supersymmetric bound states characterized by a local geometry of the form $T^{d-2k}\times\left(\mathbb{C}\times T^{2k}\right)/\mathbb{Z}_{2}^{k}$, where the $\mathbb{Z}_{2}^{k}$ act on pairs of holomorphic coordinates. In the case of an odd number of distinct reflections, we instead arrive at a non-supersymmetric configuration, namely a supersymmetric background with an additional supersymmetry breaking reflection brane added in. This geometry can be written as $T^{d-(2k+1)}\times\text{Cone}(S_{\infty}^{1}\times T^{2k}\times S^{1})/\Gamma\times\mathbb{Z}_{2}^{\text{KB}}$, where the quotient by $\mathbb{Z}_{2}^{\text{KB}}$ on the factor $S_{\infty}^{1}\times S^{1}$ is a Klein bottle (with Pin+ structure).
@@ 2.2.2 Pin+-Lifts
This story can be generalized even further by also allowing for compactifications of M-theory on non-orientable manifolds. In particular, 11D M-theory is well defined on manifolds with Pin+ structure [42, 43], which in turn suggests a further lift of $G_{U}$ to incorporate orientation reversing transformations such as reflections. [^13] From the bottom up perspective, it may appear that, for example, 9D supergravity is well-defined on Pin--manifolds. However, in order to match with the top down construction, i.e., anomaly cancellation constraints in 11D M-theory, we must restrict to the Pin+-lift.
@@
where we demand that the lifted reflection generator $\widetilde{R}$ satisfies $\widetilde{R}^{2}=1$. See Appendix C.2 for an explicit construction of the Pin+-lifts in each dimension.
@@ 4.4.1 $M>2$ Reflection Branes: Supersymmetric Case
i.e., it is an example of an $\mathfrak{so}(8)^{3}$ conformal Yukawa in the sense of reference [67]. [^32] Note in particular that these further collisions of singularities do not result in matter in even bigger representations of the flavor symmetry, but rather additional interactions. Thus, this is in accord with the conjectured absence of an isolated 5-plet of $\mathfrak{su}(2)$ discussed in [68].
@@ 4.4.2 $M>2$ Reflection Branes: Non-Supersymmetric Case
Consider next the case of configurations involving an odd number of reflection branes, say $2k+1$. In this case, we do not preserve any supersymmetries. Topologically, the geometry in this case involves a combination of the supersymmetry preserving quotient by the group $\Gamma=\mathbb{Z}_{2}^{k}$, as well as a $\mathbb{Z}_{2}^{\text{KB}}$ group action which produces the Klein bottle (KB) from a quotient of $T^{2}$, namely $\mathrm{KB}=T^{2}/\mathbb{Z}_{2}^{\text{KB}}$. In this case, the full geometry takes the form:
$T^{d-(2k+1)}\times\text{Cone}(S_{\infty}^{1}\times T^{2k}\times S^{1})/(\Gamma\times\mathbb{Z}_{2}^{\text{KB}})\text{,}$ (4.30)
where $S_{\infty}^{1}=\partial\mathbb{C}$, the circle at infinity which winds around the codimension-two defect. One way to understand these examples is to start with one of our supersymmetric backgrounds and simply add an additional non-supersymmetric reflection brane. We expect that this engineers an interacting quantum field theory.
@@ E.1 Atiyah-Hirzebruch Spectral Sequence Calculations
Definition E.9 (Wang [78, Definition 8.2]).
Let $X$ be a space and $w\in H^{2}(X;\mathbb{Z}_{2})$. An $(X,w)$-twisted spin structure on an oriented vector bundle $E\to M$ is the data of a map $f\colon M\to X$ and a trivialization of $w_{2}(E)+f^{*}(w)$. [^40] There is a more general notion of twisted spin structure allowing modifications of both $w_{1}$ and $w_{2}$: see Hebestreit-Joachim [79] as well as [80, Definition 1.24]. The James spectral sequence, and the formulas for its low-degree $d_{2}$s, both exist in this generality, as is proven in Kasprowski-Powell [81, Proposition 3.9] following Teichner [82]; see also [83, §5]. The Adams spectral sequence we use in §E.2 for twisted Spin bordism also generalizes to this setting; see [80].
@@
Theorem E.11 (Teichner [84, Proposition 1]).
@@
Teichner shows that if there is an orientable vector bundle $V\to X$ with $w_{2}(V)=w$, then the James spectral sequence is isomorphic to the Atiyah-Hirzebruch spectral sequence for the Spin bordism of the Thom spectrum $X^{V-\mathrm{rank}(V)}$ (that this is $\Omega_{*}^{\mathrm{Spin}}(X,w_{2}(V))$-twisted Spin bordism is a folk theorem, with one proof given in [23, Corollary 10.19]). It is not always possible to realize every degree-$2$ cohomology class $w$ as $w_{2}$ of a vector bundle [88, §2], and indeed this issue can occur when $w$ is the extension class for the Spin-lift of a real U-duality group $G_{U}(\mathbb{R})$ [89, Theorem 4.2], so the extra generality of the James spectral sequence is necessary. [^42] More general twisted Atiyah-Hirzebruch spectral sequences have been constructed and discussed in [90, 91, 92, 93, 94, 95, 96], but the differentials we need in the case of twisted Spin bordism have not been computed, so we use the James spectral sequence.
@@ E.2 Adams Spectral Sequence Calculations
For any space $X$ and class $w\in H^{2}(X;\mathbb{Z}_{2})$, the map $\phi\colon\Omega_{*}^{\mathrm{Spin}}(X,w)\to\Omega_{*}^{\mathrm{SO}}(X)$ is an isomorphism after localizing at any odd prime. This means that $\phi$ is an isomorphism on odd-torsion subgroups. When $w=0$, this is a standard fact (see, e.g., [23, §10.5]); for general $w$ it is less well-known but still a folklore theorem. See [100, Proof of Lemma 3.23] for a proof.
There is a map $\psi\colon\Omega_{*}^{\mathrm{SO}}(X)\to H_{*}(X;\mathbb{Z})$ obtained by sending an oriented manifold $M$ with map $f\colon M\to X$ to $f_{*}([M])$, and Thom’s computation of $\Omega_{*}^{\mathrm{SO}}$ in low degrees [101, Théorème IV.13] implies $\psi$ is an isomorphism in degrees $3$ and below. Therefore to compute the odd-primary torsion in $\Omega_{1}^{\mathrm{Spin}\text{-}\widetilde{G_{U}}}$ and $\Omega_{1}^{\mathrm{Spin}\text{-}\widetilde{G_{U}}^{+}}$, it suffices to know $H_{1}(BG_{U};\mathbb{Z})$ and $H_{1}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R}),\mathbb{Z})$, i.e. the Abelianizations of these groups, which we discussed above, e.g. in Proposition D.3. Thus in what follows we will ignore odd torsion.
@@
    Suppose you want to compute $\mathrm{Ext}_{\mathcal{A}(1)}^{s,t}(N,\mathbb{Z}_{2})$ with a minimal $\mathcal{A}(1)$-module resolution $P_{\bullet}\to N$ (see [97, §4.4]). In practice, if one only knows the structure of $N$ in degrees $d$ and below, a minimal resolution can be constructed explicitly in degrees ts<dt-s<d, simply by trying to work out the minimal resolution and stopping once higher-degree information on $N$ is necessary. This technique works because of a theoretical guarantee that, with one exception we discuss in a moment, if $N$ is concentrated in nonnegative degrees, the lowest-degree class in $P_{s}$, the $s^{\mathrm{th}}$ step of the minimal resolution, is approximately $3s$ [102, 103]. The exception is $h_{0}$-towers, which are not hard to recognize in a minimal resolution and so can be accounted for (see [97, Example 4.4.2]).
@@
    Alternatively, we can use a long exact sequence in Ext associated to a short exact sequence of $\mathcal{A}(1)$ modules, as in [97, §4.6]. Let $N$ be an $\mathcal{A}(1)$ module and $N_{>\ell}$ the submodule of $N$ consisting of elements in degrees $\ell+1$ and above. Then there is a short exact sequence
@@
Thus in particular, we may completely ignore the $\mathcal{A}(1)$-module $M$ appearing in Lemma E.18, as it cannot affect the behavior of the spectral sequence in degrees $6$ and below. [^43] In fact, because the Anderson-Brown-Peterson splitting of Spin bordism [104] generalizes to $(X,w)$-twisted Spin bordism [79, §5.2], this extends to degree $7$.
@@
$\mathrm{Sq}^{1}(1)=0\,,\enspace\mathrm{Sq}^{2}_{w}(1)=w\,,\enspace\mathrm{Sq}^{1}(r)=r^{2}\,,\enspace\mathrm{Sq}^{2}_{w}(r)=wr\,$, (E.30)
If $R_{2}$ denotes the $\mathcal{A}(1)$-module which is the kernel of the unique non-trivial $\mathcal{A}(1)$-module map $\Sigma^{-1}\mathcal{A}(1)\to\Sigma^{-1}\mathbb{Z}_{2}$ (see [97, Figure 28, left] for a picture), then in degrees $2$ and below, $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$ is isomorphic to the quotient of $R_{2}$ by its degree-$\geq 3$ elements. [^44] For the nine-dimensional U-duality group, $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$ is computed in degrees $12$ and below in [23, Proposition 14.21], and the entire $R_{2}$ summand is visible. There is also another element in degree 1, which was associated to rotations; this part is absent here, since the U-duality groups for $D\leq 7$ are perfect and the Abelianization of $(G_{U}\rtimes\mathbb{Z}_{2}^{R})$ is given by $\mathbb{Z}_{2}^{R}$ only. We can therefore write down the first few steps in a minimal resolution for $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$:

### v2

@@ 1 Introduction
Similar considerations hold for the dynamics of multiple reflection branes. We find that for reflections associated with an even number of distinct internal directions, the resulting branes form supersymmetric bound states characterized by a local geometry of the form $T^{d-2k}\times\left(\mathbb{C}\times T^{2k}\right)/\mathbb{Z}_{2}^{k}$, where the $\mathbb{Z}_{2}^{k}$ act on pairs of holomorphic coordinates. In the case of an odd number of distinct reflections, we instead arrive at a non-supersymmetric configuration, namely a supersymmetric background with an additional supersymmetry breaking reflection brane added in. This geometry can be written as $T^{d-(2k+1)}\times\text{Cone}(\mathrm{KB}_{\infty}\times T^{2k})/\Gamma$, where the $\mathrm{KB}_{\infty}$ is a circle bundle over the spacetime $S^{1}_{\infty}=\partial\mathbb{C}$ resulting in a Klein bottle (with Pin+ structure).
@@ 2.2.2 Pin+-Lifts
This story can be generalized even further by also allowing for compactifications of M-theory on non-orientable manifolds. In particular, 11D M-theory is well defined on manifolds with Pin+ structure [42, 43], which in turn suggests a further lift of $G_{U}$ to incorporate orientation reversing transformations such as reflections. [^13] In this work we are focusing on Spin-twisted duality bundles, so the overall Pin+ structure of M-theory means that we must in turn restrict to the Pin+-lift of $G_{U}$. If one considers a more general twisting by tangential structures (as can happen in some 9D supergravity backgrounds) then one might also entertain broader possible lifts. We leave the exploration of this possibility to future work.
@@
where we demand that the lifted reflection generator $\widetilde{R}$ satisfies $\widetilde{R}^{2}=1$. See Appendix C.2 for an explicit construction of the Pin+-lifts in each dimension. [^15] For $D\leq 8$ one can also define a reflection operator that squares to $(-1)^{F}$ by combining several reflections of the internal torus: see [9, §3.1].
@@ 4.4.1 $M>2$ Reflection Branes: Supersymmetric Case
i.e., it is an example of an $\mathfrak{so}(8)^{3}$ conformal Yukawa in the sense of reference [67]. [^33] Note in particular that these further collisions of singularities do not result in matter in even bigger representations of the flavor symmetry, but rather additional interactions (see e.g, [68]). Thus, this is in accord with the conjectured absence of an isolated 5-plet of $\mathfrak{su}(2)$ discussed in [69].
@@ 4.4.2 $M>2$ Reflection Branes: Non-Supersymmetric Case
Consider next the case of configurations involving an odd number of reflection branes, say $2k+1$. In this case, we do not preserve any supersymmetries. Topologically, the geometry in this case involves a combination of the supersymmetry preserving quotient by the group $\Gamma=\mathbb{Z}_{2}^{k}$, as well as a Klein bottle $\mathrm{KB}_{\infty}$, viewed as a circle bundle over the spacetime $S_{\infty}^{1}=\partial\mathbb{C}$. In this case, the full geometry takes the form:
$T^{d-(2k+1)}\times\text{Cone}(\mathrm{KB}_{\infty}\times T^{2k})/\Gamma\text{.}$ (4.30)
One way to understand these examples is to start with one of our supersymmetric backgrounds and simply add an additional non-supersymmetric reflection brane. We expect that this engineers an interacting quantum field theory.
@@ E.1 Atiyah-Hirzebruch Spectral Sequence Calculations
Definition E.9 (Wang [79, Definition 8.2]).
Let $X$ be a space and $w\in H^{2}(X;\mathbb{Z}_{2})$. An $(X,w)$-twisted spin structure on an oriented vector bundle $E\to M$ is the data of a map $f\colon M\to X$ and a trivialization of $w_{2}(E)+f^{*}(w)$. [^41] There is a more general notion of twisted spin structure allowing modifications of both $w_{1}$ and $w_{2}$: see Hebestreit-Joachim [80] as well as [81, Definition 1.24]. The James spectral sequence, and the formulas for its low-degree $d_{2}$s, both exist in this generality, as is proven in Kasprowski-Powell [82, Proposition 3.9] following Teichner [83]; see also [84, §5]. The Adams spectral sequence we use in §E.2 for twisted Spin bordism also generalizes to this setting; see [81].
@@
Theorem E.11 (Teichner [85, Proposition 1]).
@@
Teichner shows that if there is an orientable vector bundle $V\to X$ with $w_{2}(V)=w$, then the James spectral sequence is isomorphic to the Atiyah-Hirzebruch spectral sequence for the Spin bordism of the Thom spectrum $X^{V-\mathrm{rank}(V)}$ (that this is $\Omega_{*}^{\mathrm{Spin}}(X,w_{2}(V))$-twisted Spin bordism is a folk theorem, with one proof given in [23, Corollary 10.19]). It is not always possible to realize every degree-$2$ cohomology class $w$ as $w_{2}$ of a vector bundle [89, §2], and indeed this issue can occur when $w$ is the extension class for the Spin-lift of a real U-duality group $G_{U}(\mathbb{R})$ [90, Theorem 4.2], so the extra generality of the James spectral sequence is necessary. [^43] More general twisted Atiyah-Hirzebruch spectral sequences have been constructed and discussed in [91, 92, 93, 94, 95, 96, 97], but the differentials we need in the case of twisted Spin bordism have not been computed, so we use the James spectral sequence.
@@ E.2 Adams Spectral Sequence Calculations
For any space $X$ and class $w\in H^{2}(X;\mathbb{Z}_{2})$, the map $\phi\colon\Omega_{*}^{\mathrm{Spin}}(X,w)\to\Omega_{*}^{\mathrm{SO}}(X)$ is an isomorphism after localizing at any odd prime. This means that $\phi$ is an isomorphism on odd-torsion subgroups. When $w=0$, this is a standard fact (see, e.g., [23, §10.5]); for general $w$ it is less well-known but still a folklore theorem. See [101, Proof of Lemma 3.23] for a proof.
There is a map $\psi\colon\Omega_{*}^{\mathrm{SO}}(X)\to H_{*}(X;\mathbb{Z})$ obtained by sending an oriented manifold $M$ with map $f\colon M\to X$ to $f_{*}([M])$, and Thom’s computation of $\Omega_{*}^{\mathrm{SO}}$ in low degrees [102, Théorème IV.13] implies $\psi$ is an isomorphism in degrees $3$ and below. Therefore to compute the odd-primary torsion in $\Omega_{1}^{\mathrm{Spin}\text{-}\widetilde{G_{U}}}$ and $\Omega_{1}^{\mathrm{Spin}\text{-}\widetilde{G_{U}}^{+}}$, it suffices to know $H_{1}(BG_{U};\mathbb{Z})$ and $H_{1}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R}),\mathbb{Z})$, i.e. the Abelianizations of these groups, which we discussed above, e.g. in Proposition D.3. Thus in what follows we will ignore odd torsion.
@@
    Suppose you want to compute $\mathrm{Ext}_{\mathcal{A}(1)}^{s,t}(N,\mathbb{Z}_{2})$ with a minimal $\mathcal{A}(1)$-module resolution $P_{\bullet}\to N$ (see [98, §4.4]). In practice, if one only knows the structure of $N$ in degrees $d$ and below, a minimal resolution can be constructed explicitly in degrees ts<dt-s<d, simply by trying to work out the minimal resolution and stopping once higher-degree information on $N$ is necessary. This technique works because of a theoretical guarantee that, with one exception we discuss in a moment, if $N$ is concentrated in nonnegative degrees, the lowest-degree class in $P_{s}$, the $s^{\mathrm{th}}$ step of the minimal resolution, is approximately $3s$ [103, 104]. The exception is $h_{0}$-towers, which are not hard to recognize in a minimal resolution and so can be accounted for (see [98, Example 4.4.2]).
@@
    Alternatively, we can use a long exact sequence in Ext associated to a short exact sequence of $\mathcal{A}(1)$ modules, as in [98, §4.6]. Let $N$ be an $\mathcal{A}(1)$ module and $N_{>\ell}$ the submodule of $N$ consisting of elements in degrees $\ell+1$ and above. Then there is a short exact sequence
@@
Thus in particular, we may completely ignore the $\mathcal{A}(1)$-module $M$ appearing in Lemma E.18, as it cannot affect the behavior of the spectral sequence in degrees $6$ and below. [^44] In fact, because the Anderson-Brown-Peterson splitting of Spin bordism [105] generalizes to $(X,w)$-twisted Spin bordism [80, §5.2], this extends to degree $7$.
@@
$\mathrm{Sq}^{1}(1)=0\,,\kern 5.0pt\mathrm{Sq}^{2}_{w}(1)=w\,,\kern 5.0pt\mathrm{Sq}^{1}(r)=r^{2}\,,\kern 5.0pt\mathrm{Sq}^{2}_{w}(r)=wr\,$, (E.30)
If $R_{2}$ denotes the $\mathcal{A}(1)$-module which is the kernel of the unique non-trivial $\mathcal{A}(1)$-module map $\Sigma^{-1}\mathcal{A}(1)\to\Sigma^{-1}\mathbb{Z}_{2}$ (see [98, Figure 28, left] for a picture), then in degrees $2$ and below, $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$ is isomorphic to the quotient of $R_{2}$ by its degree-$\geq 3$ elements. [^45] For the nine-dimensional U-duality group, $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$ is computed in degrees $12$ and below in [23, Proposition 14.21], and the entire $R_{2}$ summand is visible. There is also another element in degree 1, which was associated to rotations; this part is absent here, since the U-duality groups for $D\leq 7$ are perfect and the Abelianization of $(G_{U}\rtimes\mathbb{Z}_{2}^{R})$ is given by $\mathbb{Z}_{2}^{R}$ only. We can therefore write down the first few steps in a minimal resolution for $H_{w}^{*}(B(G_{U}\rtimes\mathbb{Z}_{2}^{R});\mathbb{Z}_{2})$: